. The equilibrium price in a perfectly competitive market is $75. The marginal cost function is given By

MC= 4+0.2Q

What is the profit maximization rule? The firm is presently producing 40 units of output per period. To maximize profit, should the output rate be increased or decreased? Where is profit maximization show your calculations. Explain
Answer #1

Ans) In Perfectly competitive market, price is equal to marginal revenue. And a profit maximising firm in any market produces the quantity where MR = MC. (Profit maximising rule, MR = MC).

So, MR = P and to maximise profit MR = MC

75 = 4+0.2Q

75 - 4 = 0.2Q

71 = 0.2Q

Q = 71÷0.2

Q=355

Therefore, profit maximising quantity is 355 units and firm should increase the output.